given f(x)= 5x + 4 and g(x) = 3x -8 find f times g
Answer by nyc_function(2741)

f(x) = f = 5x + 4
g(x) = g = 3x - 8
f*g = (5x + 4)(3x - 8)
f*g = 15x^2 -40x + 12x - 32
f*g = 15x^2 - 28x - 32